OpenStudy (juan1857):

Find each power. Express the result in rectangular form. (1-i)^5

jhonyy9 (jhonyy9):

rewrite it in this way (1-i)^5 = (1-i)^2 *(1-i)^3 for (1-i)^2 = ? hope you know thie formula (a-b)^2 = ? for (1-i)^3 = ? hope so you know formula (a-b)^3 = ? just use these formule and will get the right answer sure but ATTENTION on i^2 = -1

OpenStudy (owikbenstan):

(-4, 4) Expand using binomial theorem, combine like terms, and find the real and imaginary parts resulting from the equation -4+4i

OpenStudy (juan1857):

pardon me i don't know the formulas

OpenStudy (juan1857):

@jhonyy9 what do you do after you break the equation

OpenStudy (freckles):

You can use binomial theorem or you can use demorive theorem

OpenStudy (freckles):

write in polar form if you want to use demorive theorem

OpenStudy (juan1857):

im just having trouble with the first steps, I know what r equal but i don't know how

OpenStudy (freckles):

\[z=a+bi \\ r=|z|=|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (juan1857):

so would it be \[\sqrt{1^{2+-i ^{2}}}\]

OpenStudy (freckles):

I don't know how you got that from what I wrote :p \[\text{ Solve } \tan(\theta)=\frac{b}{a} \text{ keeping in my that } \theta \text{ needs to be in the } \\ \text{ quadrant your point is in}\]

OpenStudy (freckles):

a=1 and b=-1 just plug them in

OpenStudy (juan1857):

how did you get -1?, sorry math isn't my thing

OpenStudy (freckles):

just by looking at 1-i

OpenStudy (freckles):

1-i is the same as 1-1i comparing this to a+bi I see a is 1 and b is -1

OpenStudy (juan1857):

ohh, now I understand, Thank You

OpenStudy (freckles):

np